# Looking Beyond Visible Cues: Implicit Video Question Answering via Dual-Clue Reasoning

### v1

@@ Abstract
Video Question Answering (VideoQA) aims to answer natural language questions based on the given video, with prior work primarily focusing on identifying the duration of relevant segments, referred to as explicit visual evidence. However, explicit visual evidence is not always directly available, particularly when questions target symbolic meanings or deeper intentions, leading to significant performance degradation. To fill this gap, we introduce a novel task and dataset, Implicit Video Question Answering (I-VQA), which focuses on answering questions in scenarios where explicit visual evidence is inaccessible. Given an implicit question and its corresponding video, I-VQA requires answering based on the contextual visual cues present within the video. To tackle I-VQA, we propose a novel reasoning framework, IRM (Implicit Reasoning Model), incorporating dual-stream modeling of contextual actions and intent clues as implicit reasoning chains. IRM comprises the Action-Intent Module (AIM) and the Visual Enhancement Module (VEM). AIM deduces and preserves question-related dual clues by generating clue candidates and performing relation deduction. VEM enhances contextual visual representation by leveraging key contextual clues. Extensive experiments validate the effectiveness of our IRM in I-VQA tasks, outperforming GPT-4o, OpenAI-o3, and fine-tuned VideoChat2 by 0.76%, 1.37%, and 4.87%, respectively. Additionally, IRM performs SOTA on similar implicit advertisement understanding and future prediction in traffic-VQA. Datasets and codes are available for double-blind review in anonymous repo: https://github.com/tychen-SJTU/Implicit-VideoQA.
@@ 1 Introduction
Video Question Answering (VideoQA) is designed to comprehend visual content and generate accurate answers by reasoning over the provided video. To facilitate reasoning through more relevant visual segments, an increasing number of VideoQA works have focused on locating key visual segments, referred to as explicit visual evidence. With the explicit evidence, reasoning models (e.g., Large Language Models) can derive more accurate answers sevila ; gqa1 ; relation . Furthermore, recent works have introduced Grounded Video Question Answering (Grounded-VQA) task nextgqa ; lita ; etbench ; rextime ; cgbench , which employs supervised training to locate explicit visual evidence more precisely.
Figure 1: Examples of Implicit VQA. Implicit VQA is more challenging due to the non-existence of explicit visual evidence. This happens when questions target symbolic meanings (left figure) or deeper intentions (right figure).
However, the performance of VideoQA is significantly compromised when explicit visual evidence is difficult to locate directly. This challenge arises from several factors, including but not limited to ambiguous visual segments ambiguous , perspective shifts perspective , and occlusions occlusions . Moreover, grounding explicit evidence becomes particularly challenging for questions that involve symbolic meanings or underlying intentions auvid ; mmau ; vvtut . In these cases, the direct identification of explicit evidence may be confounded by spurious relationships, underscoring the need for a deeper reasoning approach that goes beyond surface-level connections.
@@
To address these challenges, we set up a new task: Implict-Video Question Answring (I-VQA), which aims to infer an appropriate answer to implicit video questions, where explicit visual evidence corresponding to the given question cannot be directly accessed or located. Meanwhile, we contribute a new dataset for the training and evaluation of I-VQA by reorganizing videos from the Grounded-VQA datasets and removing the annotated visual evidence segments. However, to our knowledge, no existing solutions can directly tackle general implicit questions.
To this end, we draw inspiration from works in similar evidence-insufficient task: visual commonsense reasoning. These approaches enhance reasoning by leveraging contextual visual information and commonsense to infer additional relevant details visualcomet ; video2commonsense ; visualcommon2 . Moreover, in the video modality, richer temporal contextual inputs make it even more crucial to effectively incorporate textual information. When further considering the utilized form, cognitive science suggests Cognitive that higher-level information has greater potential for guiding reasoning. As shown in  Fig. 2, the intention conveyed in the preceding context is often reflected in subsequent actions. We can infer implicit actions by extracting intention from the preceding context, while also validating the coherence of these implicit intentions through actions in the following context. Therefore, we argue that introducing (1) context action involvement clues (2) context underlying intention clues as additional reasoning pathways is critical for implicit reasoning in the I-VQA task.
Based on this intuition, we introduce IRM, (Implicit Reasoning Model), which employs dual-stream modeling of contextual actions and intent clues as implicit reasoning chains to capture deeper underlying semantics. Specifically, we propose the Action-Intent Module (AIM), which generates textual clues representing explicit contextual actions and their underlying intentions. However, these clue candidates may be affected by factors such as hallucination MECD ; hallucination and language bias biasgpt ; biasgpt2 . To mitigate error propagation and reduce potential misleading tendencies, we incorporate a visual information verification module within AIM. This module refines dual-clue candidates by verifying them against visual information, leveraging global temporal perception for more accurate validation.
@@
Figure 2: Dual Clues Aid in Answering Implicit Video Questions. When addressing implicit problems, direct reasoning frequently fails (as indicated by the red arrows). In contrast, our proposed context clue-aided reasoning method (as indicated by the green arrows) first infers the underlying intent from the preceding context, and then verifies this intent using subsequent actions.
@@
Extensive experiments validate the effectiveness and generalizability of IRM for implicit reasoning. IRM performs SOTA in both the proposed I-VQA and current advertisement understanding datasets. Additionally, the generalization test highlights the strong performance of IRM in explicit VideoQA, particularly in predicting outcomes in SUTD-Traffic and the corresponding MVBench subset.
@@ 2 Related Work
Video Question Answering. Video question answering, as a key representative of video reasoning tasks, has seen rapid development. Recent advancements in VideoQA have been driven by the integration of pre-trained visual encoders with large language models videollama ; stllm ; vqa_1 ; vqa_2 ; vqa_3 . Notably, several works sevila ; gqa1 ; relation have focused on identifying key visual segments within the temporal domain during the reasoning process, enabling more accurate reasoning. The Grounded Video Question Answering (Grounded-VQA) task nextgqa ; lita ; etbench ; rextime ; cgbench ; timecraft ; vqa_4 ; video_thought further strengthens this approach by incorporating human annotations for supervised training. However, the performance of these methods can significantly degrade when explicit evidence is difficult to locate directly. Therefore, we propose the I-VQA task, which is specifically designed to handle the challenging implicit VideoQA.
@@
Visual Context Reasoning. Visual context reasoning refers to scenarios where the available information is insufficient for supporting the visual understanding task. Current approaches to visual context reasoning primarily focus on leveraging commonsense knowledge to infer additional visible visual information visualcomet ; video2commonsense ; visualcommon ; visualcommon2 . However, these methods are prone to reporting bias and relying on commonsense assumptions can sometimes lead to misguided conclusions. Another approach models the relationships between contexts, using known context information to infer masked events in video captioning VAR or video causal relation reasoning MECD ; mecd+ . However, these methods are susceptible to error accumulation, which can negatively affect subsequent reasoning. We offer a more robust solution, rather than explicitly inferring additional visual data, we model informative contextual clues related to action and intention. These clues form reasoning chains that capture deeper underlying semantics, while error accumulation is mitigated through a rigorous refinement process.
@@
Implicit Video Reasoning. The reasoning tasks related to implicitness in videos mainly focus on inferring implicit attributes of objects, such as mass and charge in virtual videos. ComPhy comphy requires models to identify intrinsic properties and make dynamic predictions based on these. CRIPP-VQA cripp requires physical properties to be learned from video rather than explicitly expressed in the question. ACQUIRED ACQUIRED is a combination of the two forms mentioned above. Another similar task is the PSAV adsdataset , which introduces the persuasion strategy identification task in advertisement understanding, the model needs to distinguish the implicit strategy of the advertisement video among the strategies defined by Persuasion . However, our I-VQA focuses on more general implicit questions regarding daily-life videos. Moreover, our task eliminates the possibility of taking explicit evidence as shortcuts, ensuring the rigor of implicit reasoning.
@@ 3.1 Task Definition
Moreover, to ensure the implicitness, we appropriately extend the explicit visual evidence, where the certain excluded timestamp $t_{e}^{i}$ can be represented by: $t_{e}^{i}=[max(0,t_{0}-{T}/{\sigma}),min(t,t_{1}+{T}/{\sigma})]$, where $\sigma$ indicates a constant value. Besides, following previous VideoQA tasks videollava ; qwen ; pllava ; stllm , we introduce the open-ended evaluation for a subset in the I-VQA task. After an extremely strict manual assessment, the open-ended evaluation subset is identified, which satisfies the condition of a unique correct answer without any imposed restrictions. The evaluation is represented by: $a_{o}=f_{w}(q,v)$. $a_{o}$ is the generated responses, which can be in free-form language for open-ended VideoQA.
@@ 3.2 Task Dataset
Data Source. The I-VQA task involves question-answer pairs where key visual evidence is invisible. To set up the task setting for implicit questions, we systematically reorganize and leverage explicit visual evidence annotations from widely accepted Grounded VideoQA datasets, including Next-GQA nextgqa , E.T. Bench etbench , and REXTIME rextime . For the specific implementation, we slightly extend ($\sigma$ is set to 20) and manually mask these explicit visual evidence segments in videos.
Next-GQA comprises 8,911 QA pairs and 1,557 videos from NextQA nextqa which focuses on complex temporal and causal reasoning. E.T. Bench includes 10,000 grounded QA samples, requiring models to demonstrate event-level reasoning. REXTIME consists of 12,000 QA samples sourced from ActivityNet acititynet and QV-Highlights qv-highlight and is to evaluate the ability to perform temporal reasoning.
@@
1. We exclude questions that could be correctly answered purely through commonsense reasoning. During commonsense exclusion, only questions $q$ and answer sets $\mathbb{A}$ are assessable. We exclude the examples that can be answered correctly without visual inputs $v$ (e.g., “Q: How does the boy feel after he has gotten up the first time he successfully rolled - A: Happy”). 2. We exclude questions focused on simple description questions and temporal location questions. Descriptive questions refer to questions about non-inferential details, such as the color or quantity of sth. 3. We exclude questions that are insufficient to be reasoned out with contextual visual information. Insufficient answer exclusion refers to questions that cannot be deduced from context, such as those beginning with ‘When’ asking for a time, or ‘Where’ inquiring about a location. Furthermore, we exclude questions that are manually identified as undeducible from context, detailed in the Appendix.Sec. E.
@@
These exclusions ensure that the correct answer $a^{\ast}$ of each question can be uniquely reasoned out, avoiding ambiguity in answers. Based on the three principles, we collect 5,549 QA samples, where 3,749 samples are for the training set and 1,800 samples are split into the test set.
Figure 3: Dataset Cleaning and Annotation Pipeline. We carefully check the rigor of the implicitness and filter context QA pairs with no overlap to mitigate potential information leakage. We further rewrite them as clue pairs and label the relations between acquiring clues and answering the question.
@@
Data Annotation. As discussed above, effectively leveraging clues is crucial when tackling implicit video questions. To support the following research, we have converted the QA pairs toward the same video into context clue annotations. These context QA pairs do not have overlapping durations with the target implicit question, as indicated by the blue arrows in Fig. 3. The non-overlapping condition is imposed to mitigate potential information leakage, ensuring that the clues do not directly contain parts of the answer. Specifically, we rewrite the context question into a declarative sentence as the action, and introduce the correct answer to the question as the intent as analyzed in IntentQA intentqa .
Figure 4: Dataset Statistics. The majority of videos have a duration of around one minute, with explicit evidence typically accounting for 10-30% of the video. Questions starting with ‘Why,’ ‘How,’ and ‘What’ are common. Most of the context clues transformed from context QA pairs are beneficial for answering implicit questions, with the majority containing 3-4 context clues.
@@
Moreover, to enable our work and future research to more effectively utilize these contextual clues, we annotate the relations $r=[r_{1},...,r_{n}]$ between the context pair $1\sim n$ and the implicit question $q$ using GPT-4, following the causal relation definition in REXTIME rextime and MECD MECD . Specifically, we determine whether a causal relation exists between the acquisition of context clues and the deduction of the implicit answer. Additionally, we instruct GPT-4 to pay attention to using the counterfactual inference while avoiding the illusory existence causality MECD . We label context clues contributing to implicit question answering as “1” and those that don’t as “0”. Furthermore, we manually validate the quality of 200 causal relationship annotations and find that they highly align with human expert judgment (major choice of 10 undergraduate students), with a similarity of over 92%.

### v2

@@ Abstract
Video Question Answering (VideoQA) aims to answer natural language questions based on the given video, with prior work primarily focusing on identifying the duration of relevant segments, referred to as explicit visual evidence. However, explicit visual evidence is not always directly available, particularly when questions target symbolic meanings or deeper intentions, leading to significant performance degradation. To fill this gap, we introduce a novel task and dataset, Implicit Video Question Answering (I-VQA), which focuses on answering questions in scenarios where explicit visual evidence is inaccessible. Given an implicit question and its corresponding video, I-VQA requires answering based on the contextual visual cues present within the video. To tackle I-VQA, we propose a novel reasoning framework, IRM (Implicit Reasoning Model), incorporating dual-stream modeling of contextual actions and intent clues as implicit reasoning chains. IRM comprises the Action-Intent Module (AIM) and the Visual Enhancement Module (VEM). AIM deduces and preserves question-related dual clues by generating clue candidates and performing relation deduction. VEM enhances contextual visual representation by leveraging key contextual clues. Extensive experiments validate the effectiveness of our IRM in I-VQA tasks, outperforming GPT-4o, OpenAI-o3, and fine-tuned VideoChat2 by 0.76%, 1.37%, and 4.87%, respectively. Additionally, IRM performs SOTA on similar implicit advertisement understanding and future prediction in traffic-VQA. Datasets and codes are available in the repo: https://github.com/tychen-SJTU/Implicit-VideoQA.
keywords:
Video Question Answering, Video Understanding, Video Reasoning, Implicit Reasoning
Figure 1: Examples of Implicit VQA. Implicit VQA is more challenging due to the non-existence of explicit visual evidence. This happens when questions target symbolic meanings (left figure) or deeper intentions (right figure). Although some information in the question can be directly extracted from the video, and the behavioral action information can be easily captured (as shown by the green arrows), implicit intentions or inner emotions cannot be simply or directly inferred (as shown by the red arrows).
@@ 1 Introduction
Video Question Answering (VideoQA) is designed to comprehend visual content and generate accurate answers by reasoning over the provided video. To facilitate reasoning through more relevant visual segments, an increasing number of VideoQA works have focused on locating key visual segments, referred to as explicit visual evidence. With the explicit evidence, reasoning models (e.g., Large Language Models) can derive more accurate answers [sevila, gqa1, relation]. Furthermore, recent works have introduced Grounded Video Question Answering (Grounded-VQA) task [nextgqa, lita, etbench, rextime, cgbench], which employs supervised training to locate explicit visual evidence more precisely.
However, the performance of VideoQA is significantly compromised when explicit visual evidence is difficult to locate directly. This challenge arises from several factors, including but not limited to ambiguous visual segments [ambiguous], perspective shifts [perspective], and occlusions [occlusions]. Moreover, grounding explicit evidence becomes particularly challenging for questions that involve symbolic meanings or underlying intentions [auvid, mmau, vvtut]. In these cases, the direct identification of explicit evidence may be confounded by spurious relationships, underscoring the need for a deeper reasoning approach that goes beyond surface-level connections.
@@
To address these challenges, we set up a new task: Implict-Video Question Answring (I-VQA), which aims to infer an appropriate answer to implicit video questions, where explicit visual evidence corresponding to the given question cannot be directly accessed or located. Meanwhile, we contribute a new dataset for the training and evaluation of I-VQA: We construct and validate implicit video question answering through a semi-automated process, based on existing Grounded-VQA datasets. However, to our knowledge, no existing solutions can directly tackle implicit questions.
Figure 2: Dual Clues Aid in Answering Implicit Video Questions. When facing implicit questions, direct reasoning fails (as indicated by the red arrows). In contrast, we proposed an effective contextual clue-aided reasoning method (as indicated by the green arrows). Contextual reasoning paths are constructed by first inferring underlying intent from the preceding context, and then verifying through subsequent actions.
To this end, we draw inspiration from works in similar evidence-insufficient task: visual commonsense reasoning. These approaches enhance reasoning by leveraging contextual visual information and commonsense to infer additional relevant details [visualcomet, video2commonsense, visualcommon2, informative, expressive]. Moreover, in the video modality, richer temporal contextual inputs make it even more crucial to effectively incorporate contextual information. When further considering the utilized form, cognitive science suggests [Cognitive] that higher-level information has greater potential for guiding reasoning. As shown in  Fig. 2, the intention conveyed in the preceding context is often reflected in subsequent actions. We can infer implicit actions by extracting intention from the preceding context, while also validating the coherence of these implicit intentions through actions in the following context. Therefore, we argue that introducing dual clues: (1) context action involvement clues (2) context underlying intention clues as additional reasoning pathways is critical for implicit reasoning in the I-VQA task.
Based on this intuition, we introduce IRM, (Implicit Reasoning Model), which employs dual-stream modeling of contextual actions and intent clues as implicit reasoning chains to capture deeper underlying semantics. Specifically, we propose the Action-Intent Module (AIM), which deduces textual clue candidates representing explicit contextual actions and their underlying intentions. However, these clue candidates may be affected by factors such as hallucination [MECD, hallucination] and language bias [biasgpt, biasgpt2]. To mitigate error propagation and reduce potential misleading tendencies, we incorporate a visual information verification module within AIM. This module refines dual-clue candidates by verifying them against visual information, leveraging global temporal perception for more accurate validation.
@@
Evaluation of I-VQA using existing powerful multimodal understanding models, such as GPT-4o, shows that their multi-choice performance is only slightly above 50%, significantly below human-level performance and the results achieved on other widely-used video datasets. This underscores the considerable room for improvement in implicit video question answering. Furthermore, extensive experiments demonstrate the effectiveness and generalizability of IRM for implicit reasoning. IRM achieves state-of-the-art performance not only on the proposed I-VQA dataset but also on current advertisement understanding datasets. Additionally, generalization tests highlight IRM’s robust performance in general VideoQA tasks, particularly in predicting outcome events in the SUTD-Traffic dataset and performing general reasoning in MVBench.
  • We introduced the paradigm of implicit video question answering, highlighted the limitations of current models in implicit reasoning capabilities, and emphasized its importance as a largely overlooked yet crucial research direction within VideoQA.
  • We developed the first dataset for implicit video question answering to our knowledge, named I-VQA, through a carefully designed semi-automated pipeline.
  • We demonstrated a robust method for performing implicit reasoning—leveraging dual-stream contextual clue reasoning—and validated its generalizability, showing that it does not compromise the model’s performance on explicit reasoning tasks.
@@ 2 Related Work
Video question answering, as a key representative of video reasoning tasks, has seen rapid development. Recent advancements in VideoQA have been driven by the integration of pre-trained visual encoders with large language models [videollama, stllm, vqa_1, vqa_2, vqa_3]. Notably, several works [sevila, gqa1, relation] have focused on identifying key visual segments within the temporal domain during the reasoning process, enabling more accurate reasoning. The Grounded Video Question Answering (Grounded-VQA) task [nextgqa, lita, etbench, rextime, cgbench, timecraft, vqa_4, video_thought] further strengthens this approach by incorporating human annotations for supervised training. However, the performance of these methods can significantly degrade when explicit evidence is difficult to locate directly. Therefore, we propose the I-VQA task, which is specifically designed to handle the challenging implicit VideoQA.
2.2 Visual Context Reasoning.
@@
Visual context reasoning refers to scenarios where the available information is insufficient for supporting the visual understanding task. Current approaches to visual context reasoning primarily focus on leveraging commonsense knowledge to infer additional visible visual information [visualcomet, video2commonsense, visualcommon, visualcommon2]. However, these methods are prone to reporting bias and relying on commonsense assumptions can sometimes lead to misguided conclusions. Another approach models the relationships between contexts, using known context information to infer masked events in video captioning [VAR] or video causal relation reasoning [MECD, mecd+]. However, these methods are susceptible to error accumulation, which can negatively affect subsequent reasoning. We offer a more robust solution to tackle targeted implicit questions; rather than explicitly inferring additional visual data, we construct informative contextual clues consisting of action involvement and semantic intention. These clues form reasoning chains that capture deeper underlying semantics, while error accumulation is mitigated through a rigorous refinement process.
2.3 Implicit Video Reasoning.
@@
The reasoning tasks related to implicitness in videos mainly focus on inferring implicit attributes of objects, such as mass and charge in virtual videos. ComPhy [comphy] requires models to identify intrinsic properties and make dynamic predictions based on these. CRIPP-VQA [cripp] requires physical properties to be learned from video rather than explicitly expressed in the question. ACQUIRED [ACQUIRED] is a combination of the two forms mentioned above. Another similar task is the PSAV [adsdataset], which introduces the persuasion strategy identification task in advertisement understanding, the model needs to distinguish the implicit strategy of the advertisement video among the strategies defined by [Persuasion]. However, by constructing a new dataset, our I-VQA focuses on more general implicit questions regarding daily-life videos. Moreover, our I-VQA eliminates the possibility of taking explicit evidence as shortcuts, ensuring the rigor of implicit reasoning. Therefore, it enables a more rigorous and comprehensive evaluation of the model’s implicit reasoning capability.
@@ 3 I-VQA Task
Figure 3: Dataset Cleaning and Annotation Pipeline. We carefully check the rigor of the question’s implicitness through a semi-automated pipeline, and filter context QA pairs with no overlap to mitigate potential information leakage. We further rewrite context QA pairs as action-intent clues and label the relations between acquiring clues and answering the question.
@@ 3.1 Task Definition
Moreover, to ensure the implicitness, we appropriately extend the explicit visual evidence, where the certain excluded timestamp $t_{e}^{i}$ can be represented by: $t_{e}^{i}=[max(0,t_{0}-{T}/{\sigma}),min(t,t_{1}+{T}/{\sigma})]$, where $\sigma$ indicates a constant value. Besides, following previous VideoQA tasks [videollava, qwen, pllava, stllm], we introduce the open-ended evaluation for a subset in the I-VQA task. After an extremely strict manual assessment, the open-ended evaluation subset is identified, which satisfies the condition of a unique correct answer without any imposed restrictions. The evaluation is represented by: $a_{o}=f_{w}(q,v)$. $a_{o}$ is the generated responses, which can be in free-form language for open-ended VideoQA.
Figure 4: Dataset Statistics. The majority of videos have a duration of around one minute, with explicit evidence typically accounting for 10-30% of the video. Questions starting with ‘Why’, ‘How’, and ‘What’ are common. Most of the context clues transformed from context QA pairs are beneficial for answering implicit questions, with the majority containing 3-4 context clues.
@@ 3.2 Task Dataset
Data Source. The I-VQA task involves question-answer pairs where key visual evidence is invisible. To set up the task setting for implicit questions, we systematically reorganize and leverage explicit visual evidence annotations from widely accepted Grounded VideoQA datasets, including Next-GQA [nextgqa], E.T. Bench [etbench], and REXTIME [rextime]. For the specific implementation, we slightly extend ($\sigma$ is set to 20) and manually mask these explicit visual evidence segments in videos.
Next-GQA comprises 8,911 QA pairs and 1,557 videos from NextQA [nextqa] which focuses on complex temporal and causal reasoning. E.T. Bench includes 10,000 grounded QA samples, requiring models to demonstrate event-level reasoning. REXTIME consists of 12,000 QA samples sourced from ActivityNet [acititynet] and QV-Highlights [qv-highlight] and is to evaluate the ability to perform temporal reasoning.
@@
1. We exclude questions that could be correctly answered purely through commonsense reasoning. During the commonsense exclusion phase, only the questions, denoted as $q$, and their corresponding answer sets, $\mathbb{A}$, were provided for assessment. The primary objective was to eliminate instances that could be answered correctly without the associated visual input $v$. A quintessential example of such an excluded question is, “Q: How does the boy feel after he has gotten up the first time he successfully rolled? – A: Happy.” To do this, five annotators identified questions answerable by commonsense alone, and the Pearson Correlation of inter-annotator agreement achieves 0.95 (ranging from -1 to 1, more relative when closer to 1). Besides, our protocol was stringent: a question was removed if even one annotator flagged it. Finally, the effectiveness of this process was validated when GPT-4o, operating without visual input, performed at a level close to random chance on the test set, confirming the rigor of our screening [mecd+, egoschema].
2. We exclude questions focused on simple description questions and temporal location questions. Descriptive questions refer to questions about non-inferential details, such as the color or quantity of something. The screening for these questions was conducted by GPT-4 on the input questions, and the rigor of this process was further ensured by manual review. Specifically, GPT-4 removed around 2,000 examples from 10,000. We verified that these removals met the criteria. Then, we conducted a further review of the remaining examples using cross-validation by five annotators. The Pearson Correlation between the annotators reached nearly 0.98. Consistent cases were excluded, and a meeting was held to decide on the removal of controversial examples, resulting in an additional 168 exclusions in total. Temporal location questions, on the other hand, were filtered out through a string-matching method to remove QA pairs whose answers contained specific timestamps.
@@
3. We exclude questions that are insufficient to be reasoned out with contextual visual information. Insufficient answer exclusion refers to questions that cannot be deduced from context, such as those beginning with ‘When’ asking for a time, or ‘Where’ inquiring about a location. This exclusion was also implemented by a string-matching method. Furthermore, we exclude questions that are manually identified as undeducible from context through a similar rigorous human examination; a detailed case is shown in the Appendix.Sec. E.
@@
These exclusions ensure that the correct answer $a^{\ast}$ of each question can be uniquely reasoned out, avoiding ambiguity in answers. Based on the three principles, we collect 5,549 QA samples from about 30,000 original QA samples, where 3,749 samples of I-VQA are for the training set and 1,800 samples are split into the test set.
Data Annotation. As discussed above, effectively leveraging clues is crucial when tackling implicit video questions. To support future research, we have annotated contextual clues for all implicit questions in the I-VQA training set.
@@
The annotation process consists of two main steps. First, we generate numerous “action-intent” pairs. To do this, we rewrite the context question into a declarative sentence to serve as the “action,” and use the correct answer to the question as the “intent,” following the analysis in IntentQA [intentqa]. As indicated by the blue arrows in Fig. 3, these contextual QA pairs have time durations that do not overlap with the target implicit question. This non-overlapping condition is imposed to mitigate potential information leakage, ensuring that the clues do not directly contain parts of the answer.
@@
Next, we filter the generated “action-intent” pairs to select clues that are relevant to the implicit question. We annotate the relations $r=[r_{1},...,r_{n}]$ between the context pair $1\sim n$ and the implicit question $q$ using GPT-4, following the causal relation definition in REXTIME [rextime] and MECD [MECD]. Specifically, we determine whether a causal relation exists between the acquisition of context clues and the deduction of the implicit answer. Additionally, we instruct GPT-4 to pay attention to using the counterfactual inference while avoiding the illusory existence causality [MECD]. We label context clues contributing to implicit question answering as “1” and those that don’t as “0”. Furthermore, we manually validate the quality of 200 causal relationship annotations and find that they highly align with human expert judgment (major choice of 10 annotators with video understanding research experiments), with a similarity of over 92% (which is robust enough for model learning to distinguish which clues contribute to implicit reasoning [MECD, rextime]).